# Anomalous Coherence Length in Superconductors with Quantum Metric

### v5

@@ II.1 Proof of the independence of coherence length from orbital positions
$C(\bm{r}_{ij})$ $=\sum_{\alpha}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle,$ (S29)
with $\bm{r}_{ij}=\bm{r}_{i}-\bm{r}_{j}$. The correlator $C(\bm{r}_{ij})$ is expected to be decay exponentially $C(\bm{r}_{ij})\sim e^{-|\bm{r}_{ij}|/\xi}$ in an isotropic system. We aim to demonstrate that the correlator in Eq. (S29) is invariant under the choice of $\bm{\delta}_{\alpha}$. We reformulate the mean-field Hamiltonian in the band basis of $H_{0}$. We can expand the orbital fermions in the band basis
$a_{i\alpha\sigma}^{\dagger}=\frac{1}{\sqrt{N}}\sum_{n\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}u_{n\alpha\sigma}(\bm{k})c_{n\sigma}^{\dagger}(\bm{k}),$ (S30)
@@
$\sum_{\beta}h_{\alpha\beta}^{\sigma}(\bm{k})u_{n\beta\sigma}(\bm{k})=\epsilon_{n}(\bm{k})u_{n\alpha\sigma}(\bm{k}),$ (S31)
@@
$H=\sum_{n\bm{k}\sigma}\epsilon_{n}(\bm{k})c_{n\bm{k}\sigma}^{\dagger}c_{n\bm{k}\sigma}+\sum_{\bm{k}}\sum_{nm}\left[c_{n\uparrow}^{\dagger}(\bm{k})c_{m\downarrow}^{\dagger}(-\bm{k})\Delta_{nm}+h.c.\right],$ (S32)
@@
$C(\bm{r}_{ij})$ $=\sum_{\alpha}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle$
$=\frac{T}{N}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}_{i}-\bm{r}_{j})}\sum_{nmhd}\sum_{\alpha\bm{k}}u_{n\alpha}^{*}(\bm{k}+\bm{q})u_{m\alpha}(\bm{k})u_{h\alpha}^{*}(\bm{k}^{\prime})u_{d\alpha}(\bm{k}^{\prime}+\bm{q})\langle c_{n\downarrow}(\bm{k}+\bm{q})c_{m\uparrow}(-\bm{k})c_{h\uparrow}^{\dagger}(-\bm{k}^{\prime})c_{d\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle.$ (S33)
@@
$h_{\alpha\beta}^{\sigma}(\bm{k})$ $\rightarrow h_{\alpha\beta}^{\sigma}(\bm{k})e^{-i\bm{k}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})},$ (S34)
$u_{\alpha}(\bm{k})$ $\rightarrow e^{i\bm{k}\cdot\bm{x}_{\alpha}}u_{\alpha}(\bm{k}).$ (S35)
It is easy to check that $u_{n\alpha}^{*}(\bm{k}+\bm{q})u_{m\alpha}(\bm{k})u_{h\alpha}^{*}(\bm{k}^{\prime})u_{d\alpha}(\bm{k}^{\prime}+\bm{q})$ and $\Delta_{nm}$ are invariant under the choice of $\bm{\delta}_{\alpha}$. Moreover, the correlator $\langle c_{n\downarrow}(\bm{k}+\bm{q})c_{m\uparrow}(-\bm{k})c_{h\uparrow}^{\dagger}(-\bm{k}^{\prime})c_{d\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle$ in Eq. (II.1) can be obtained based on the mean-field Hamiltonian in Eq. (S32), which is also invariant under the choice of $\bm{\delta}_{\alpha}$. Therefore, we can conclude that the Cooper correlator $C(\bm{r}_{ij})$ is variant under the choice of $\bm{\delta}_{\alpha}$, which directly implies that the coherence length is a quantity which is independent of the intra-cell-orbital positions.
II.2 Uniform pairing condition and minimum quantum metric
@@
$\mathcal{M}(\bm{q})$ $=\sum_{ij\alpha}e^{-i\bm{q}\cdot\bm{r}_{ij}}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle$
$=\frac{1}{N}\sum_{\bm{k}\bm{k}^{\prime}}\sum_{\alpha}\langle a_{\alpha\downarrow}(\bm{k}+\bm{q})a_{\alpha\uparrow}(-\bm{k})a_{\alpha\uparrow}^{\dagger}(-\bm{k}^{\prime})a_{\alpha\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle$
$\rightarrow\frac{T}{N}\sum_{\bm{k}}\sum_{\alpha}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|^{2}\mathrm{Tr}G(i\omega_{n},\bm{k}+\bm{q})G(-i\omega_{n},-\bm{k}).$ (S36)
where $G$ is the Gorkov Green function in Eq. (S20). Obviously, $\mathcal{M}(\bm{q})$ remains invariant under the choice of $\bm{\delta}_{\alpha}$ since both $|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|^{2}$ and $\langle c_{\downarrow}(\bm{k}+\bm{q})c_{\uparrow}(-\bm{k})c_{\uparrow}^{\dagger}(-\bm{k})c_{\downarrow}^{\dagger}(\bm{k}+\bm{q})\rangle$ are invariant. The quantity $\mathcal{M}(\bm{q})$ is then expected to form as
$\mathcal{M}(\bm{q})=\frac{\mathcal{M}(0)}{1+q^{2}\xi^{2}}$ (S37)
@@
$\xi^{2}=-\frac{1}{2\mathcal{M}(\bm{0})}\frac{\partial^{2}\mathcal{M}(\bm{q})}{\partial q^{2}}|_{q=0}$ (S38)
@@
$=\frac{1}{N_{orb}}|\langle u(\bm{k}+\bm{q})|u(\bm{k})\rangle|^{2},$ (S39)
where $N_{orb}$ is the number of orbitals. There are two aspects of this treatment. First, it assumes that the $N_{orb}$ orbitals fulfill a condition that $1/N\sum_{\bm{k}}|u_{\alpha}(\bm{k})|^{2}=1/N_{orb}$, which leads to the same pairing gap of all orbitals. In Refs. Huhtinen et al. (2022b); torma2018quantum, this condition is referred to as the uniform pairing condition. Second, however, such assumption in generally fails to keep the invariance under the choice of the intra-unit cell orbital positions. Under the shift $\bm{\delta}_{\alpha}\rightarrow\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$, and $u_{\alpha}(\bm{k})\rightarrow u_{\alpha}(\bm{k})e^{i\bm{k}\cdot\bm{x}_{\alpha}}$, one can find that
@@
$=$ $\sum_{\alpha\beta}e^{-i\bm{q}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|e^{i\theta_{\bm{k}\alpha}(\bm{q})}|u_{\beta}(\bm{k}+\bm{q})u_{\beta}^{*}(\bm{k})|e^{-i\theta_{\bm{k}\beta}(\bm{q})},$ (S40)
This may lead to an unphysical contribution. In the expression before the projection, the quantity does not depend on the phase factor $\theta_{\alpha}(\bm{q})$. When the phase factor vanishes, the assumption in Eq.(II.2) is valid. For it, we can shift the orbital positions by $\bm{\delta}_{\alpha}\rightarrow\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$, then
$\exp\left[-i(\theta_{\bm{k}\alpha}(\bm{q})-\theta_{\bm{k}\beta}(\bm{q}))\right]\rightarrow\exp\left[-i(\theta_{\bm{k}\alpha}(\bm{q})-\theta_{\bm{k}\beta}(\bm{q}))\right]e^{-i\bm{q}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})},$ (S41)
@@
$\frac{1}{N}\sum_{\bm{k}}\frac{d}{dq_{a}}\theta_{\bm{k}\alpha\beta}(\bm{q})|_{\bm{q}\rightarrow 0}=\bm{x}_{\alpha}^{a}-\bm{x}_{\beta}^{a}.$ (S42)
In this case, we will find that the quantum metric indeed becomes the minimal one. We can fix the position of the first orbital $\bm{\delta}_{1}=0$ and $\bm{x}_{1}=0$. The shift of other orbitals can be accordingly determined by Eq. (S42). Furthermore, due to the dependence of $\theta_{\bm{k}\alpha\beta}(\bm{q})$ on the momentum $\bm{k}$, in general it is impossible to choose a set of orbital positions such that the second term in Eq. (II.2) vanishes. This is different from the case of superfluid weight. In fact, the best we can do is to reach the minimal quantum metric in Eq. (II.2). In Ref.Huhtinen et al. (2022b), it was shown that the resulting $\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$ for the minimal quantum metric is unique and independent of the initial positions $\bm{\delta}_{\alpha}$. Therefore, we can reach a conclusion that the coherence length $\xi$ for a superconductor is controlled by the minimal quantum metric under the uniform pairing condition.
One may find that the coherence length obtained under the uniform pairing condition is slightly larger than the value by directly evaluating the $\mathcal{M}(\bm{q})$ in Eq. (II.2). There are two origins of the errors. Suppose that we start with the orbital positions corresponding to the minimal quantum metric. The first type of the error comes from the factor $\mathcal{M}(0)$ in Eq. (S38), and the error can be evaluated from the Chebyshev’s sum inequality,
$\frac{1}{N}\sum_{\bm{k}\alpha}|u_{\alpha}(\bm{k})|^{4}-\frac{1}{NN_{orb}}\sum_{\bm{k}}(\sum_{\alpha}|u_{\alpha}(\bm{k})|^{2})^{2}=\frac{1}{NN_{orb}}\sum_{\bm{k},\alpha\beta}(|u_{\alpha}(\bm{k})|^{2}-|u_{\beta}(\bm{k})|^{2})^{2}\geq 0.$ (S43)
@@
$\sigma^{2}_{ab}=\frac{1}{2N}\sum_{\bm{k}\alpha\beta}(\partial_{a}\theta_{\bm{k}\alpha\beta}(0)-\overline{\partial_{a}\theta_{\alpha\beta}(0)})(\partial_{b}\theta_{\bm{k}\alpha\beta}(0)-\overline{\partial_{b}\theta_{\alpha\beta}(0)}),$ (S44)
@@
III Coherence length for a weakly dispersive band
III.1 Derivation on the total coherence length
@@
$|\Lambda(\bm{k}+\bm{q},\bm{k})|^{2}$ $=1-\sum_{ab}\mathcal{G}_{ab}(\bm{k})q_{a}q_{b}+\mathcal{O}(q^{3}),$ (S45)
@@
$\mathcal{G}_{ab}(\bm{k})=\mathrm{Re}\left[\langle\partial_{a}u(\bm{k})|\partial_{b}u(\bm{k})\rangle-\langle\partial_{a}u(\bm{k})|u(\bm{k})\rangle\langle u(\bm{k})|\partial_{b}u(\bm{k})\rangle\right],$ (S46)
Since the Green function in Eq. (II.1) is independent on the momentum in the flatband limit, we can obtain the tendency
@@
$=\frac{\tanh\left(\frac{\beta\sqrt{\Delta^{2}+\mu^{2}}}{2}\right)}{\sqrt{\Delta^{2}+\mu^{2}}}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\frac{1}{1+\sum_{ab}\overline{\mathcal{G}}_{ab}q_{a}q_{b}},$ (S47)
@@
$=$ $T\sum_{n}\frac{2}{\omega_{n}^{2}+\mu{}^{2}+\Delta^{2}}=\frac{\tanh\left(\frac{\beta\sqrt{\Delta^{2}+\mu^{2}}}{2}\right)}{\sqrt{\Delta^{2}+\mu^{2}}},$ (S48)
@@
$=\frac{\tanh\left(\frac{\beta}{2}\varepsilon\right)}{\varepsilon\sqrt{\det\bar{\mathcal{G}}_{ab}}}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}}{q^{2}+\xi^{-2}},$ (S49)
@@
$\xi=(\det\overline{\mathcal{G}}_{ab})^{\frac{1}{4}}.$ (S50)
@@
$\mathcal{C}(\bm{r},\bm{r}^{\prime})\equiv\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\mathcal{M}(\bm{q}),$ (S51)
@@
$\mathcal{M}(\bm{q})=\frac{T}{N}\sum_{n\bm{k}}|\Lambda(\bm{k}+\bm{q},\bm{k})|^{2}\mathrm{Tr}[G(i\omega_{n},\bm{k}+\bm{q})G(-i\omega_{n},-\bm{k})].$ (S52)
@@
$\mathcal{M}(\bm{q})=\mathcal{M}_{0}+\mathcal{M}_{\mathrm{qm}}(\bm{q})+\mathcal{M}_{\mathrm{con}}(\bm{q}).$ (S53)
@@
$\equiv\mathcal{M}_{0}(\Delta,\mu),$ (S54)
@@
$\nu=\int\frac{d^{2}k}{(2\pi)^{2}}\theta(\mu_{0}-\epsilon(\bm{k})),$ (S55)
@@
$\rho_{0}(\epsilon_{F})=\frac{1}{W}.$ (S56)
Then we can calculate $\mathcal{M}_{0}(\Delta,\mu)$ in Eq. (S54) at zero temperature,
@@
$=\frac{1}{\sqrt{\Delta^{2}+\mu^{2}}}.$ (S57)
@@
$=-q^{2}\mathcal{M}_{0}(\Delta,\mu)\sqrt{\det\overline{\mathcal{G}}_{ab}},$ (S58)
@@
$\overline{\mathcal{G}}_{ab}=\frac{\frac{1}{N}\sum_{\bm{k}}\mathcal{G}_{ab}(\bm{k})\frac{1}{\varepsilon(\bm{k})}}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\mathcal{G}_{ab}(\bm{k})}{\varepsilon(\bm{k})}}{\mathcal{M}_{0}(\Delta,\mu)},$ (S59)
@@
$=-\frac{T}{N}\sum_{n\bm{k}}\frac{2\bm{v}_{F}\cdot\bm{q}[\epsilon(\bm{k})-\mu+\bm{v}_{F}\cdot\bm{q}]}{[\Delta^{2}+(\epsilon(\bm{k})-\mu)^{2}+\omega_{n}^{2}][\Delta^{2}+(\epsilon(\bm{k})-\mu+\bm{v}_{F}\cdot\bm{q})^{2}+\omega_{n}^{2}]},$ (S60)

### v6

@@ VIII Competing Interests
$C(\bm{r}_{ij})$ $=\sum_{\alpha}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle_{c}$ (29)
$\equiv\sum_{\alpha}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle-\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}\rangle\langle a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle,$ (30)
with $\bm{r}_{ij}=\bm{r}_{i}-\bm{r}_{j}$. The correlator $C(\bm{r}_{ij})$ is expected to be decay exponentially $C(\bm{r}_{ij})\sim e^{-|\bm{r}_{ij}|/\xi}$ in an isotropic system. We aim to demonstrate that the correlator in Eq. (30) is invariant under the choice of $\bm{\delta}_{\alpha}$. We reformulate the mean-field Hamiltonian in the band basis of $H_{0}$. We can expand the orbital fermions in the band basis
$a_{i\alpha\sigma}^{\dagger}=\frac{1}{\sqrt{N}}\sum_{n\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}u_{n\alpha\sigma}(\bm{k})c_{n\sigma}^{\dagger}(\bm{k}),$ (31)
@@
$\sum_{\beta}h_{\alpha\beta}^{\sigma}(\bm{k})u_{n\beta\sigma}(\bm{k})=\epsilon_{n}(\bm{k})u_{n\alpha\sigma}(\bm{k}),$ (32)
@@
$H=\sum_{n\bm{k}\sigma}\epsilon_{n}(\bm{k})c_{n\bm{k}\sigma}^{\dagger}c_{n\bm{k}\sigma}+\sum_{\bm{k}}\sum_{nm}\left[c_{n\uparrow}^{\dagger}(\bm{k})c_{m\downarrow}^{\dagger}(-\bm{k})\Delta_{nm}+h.c.\right],$ (33)
@@
$C(\bm{r}_{ij})$ $=\sum_{\alpha}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle_{c}$
$=\frac{T}{N}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}_{i}-\bm{r}_{j})}\sum_{nmhd}\sum_{\alpha\bm{k}}u_{n\alpha}^{*}(\bm{k}+\bm{q})u_{m\alpha}(\bm{k})u_{h\alpha}^{*}(\bm{k}^{\prime})u_{d\alpha}(\bm{k}^{\prime}+\bm{q})\langle c_{n\downarrow}(\bm{k}+\bm{q})c_{m\uparrow}(-\bm{k})c_{h\uparrow}^{\dagger}(-\bm{k}^{\prime})c_{d\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle_{c}.$ (34)
@@
$h_{\alpha\beta}^{\sigma}(\bm{k})$ $\rightarrow h_{\alpha\beta}^{\sigma}(\bm{k})e^{-i\bm{k}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})},$ (35)
$u_{\alpha}(\bm{k})$ $\rightarrow e^{i\bm{k}\cdot\bm{x}_{\alpha}}u_{\alpha}(\bm{k}).$ (36)
It is easy to check that $u_{n\alpha}^{*}(\bm{k}+\bm{q})u_{m\alpha}(\bm{k})u_{h\alpha}^{*}(\bm{k}^{\prime})u_{d\alpha}(\bm{k}^{\prime}+\bm{q})$ and $\Delta_{nm}$ are invariant under the choice of $\bm{\delta}_{\alpha}$. Moreover, the correlator $\langle c_{n\downarrow}(\bm{k}+\bm{q})c_{m\uparrow}(-\bm{k})c_{h\uparrow}^{\dagger}(-\bm{k}^{\prime})c_{d\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle_{c}$ in Eq. (Proof of the independence of coherence length from orbital positions) can be obtained based on the mean-field Hamiltonian in Eq. (33), which is also invariant under the choice of $\bm{\delta}_{\alpha}$. Therefore, we can conclude that the Cooper correlator $C(\bm{r}_{ij})$ is variant under the choice of $\bm{\delta}_{\alpha}$, which directly implies that the coherence length is a quantity which is independent of the intra-cell-orbital positions.
Uniform pairing condition and minimum quantum metric
@@
$\mathcal{M}(\bm{q})$ $=\sum_{ij\alpha}e^{-i\bm{q}\cdot\bm{r}_{ij}}\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}a_{j\alpha\uparrow}^{\dagger}a_{j\alpha\downarrow}^{\dagger}\rangle_{c}$
$=\frac{1}{N}\sum_{\bm{k}\bm{k}^{\prime}}\sum_{\alpha}\langle a_{\alpha\downarrow}(\bm{k}+\bm{q})a_{\alpha\uparrow}(-\bm{k})a_{\alpha\uparrow}^{\dagger}(-\bm{k}^{\prime})a_{\alpha\downarrow}^{\dagger}(\bm{k}^{\prime}+\bm{q})\rangle_{c}$
$\rightarrow\frac{T}{N}\sum_{\bm{k}}\sum_{\alpha}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|^{2}G_{0}(i\omega_{n},\bm{k}+\bm{q})G_{0}(-i\omega_{n},-\bm{k}).$ (37)
where $G_{0}$ is the normal Gorkov’s Green function in Eq. (20). Obviously, $\mathcal{M}(\bm{q})$ remains invariant under the choice of $\bm{\delta}_{\alpha}$ since both $|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|^{2}$ and $\langle c_{\downarrow}(\bm{k}+\bm{q})c_{\uparrow}(-\bm{k})c_{\uparrow}^{\dagger}(-\bm{k})c_{\downarrow}^{\dagger}(\bm{k}+\bm{q})\rangle_{c}$ are invariant. The quantity $\mathcal{M}(\bm{q})$ is then expected to form as
$\mathcal{M}(\bm{q})=\frac{\mathcal{M}(0)}{1+q^{2}\xi^{2}}~{},$ (38)
@@
$\xi^{2}=-\frac{1}{2\mathcal{M}(\bm{0})}\frac{\partial^{2}\mathcal{M}(\bm{q})}{\partial q^{2}}|_{q=0}~{}.$ (39)
@@
$=\frac{1}{N_{orb}}|\langle u(\bm{k}+\bm{q})|u(\bm{k})\rangle|^{2},$ (40)
where $N_{orb}$ is the number of orbitals. There are two aspects of this treatment. First, it assumes that the $N_{orb}$ orbitals fulfill a condition that $1/N\sum_{\bm{k}}|u_{\alpha}(\bm{k})|^{2}=1/N_{orb}$, which leads to the same pairing gap of all orbitals. In Refs. Huhtinen et al. (2022); Törmä et al. (2018), this condition is referred to as the uniform pairing condition. Second, however, such assumption in generally fails to keep the invariance under the choice of the intra-unit cell orbital positions. Under the shift $\bm{\delta}_{\alpha}\rightarrow\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$, and $u_{\alpha}(\bm{k})\rightarrow u_{\alpha}(\bm{k})e^{i\bm{k}\cdot\bm{x}_{\alpha}}$, one can find that
@@
$=$ $\sum_{\alpha\beta}e^{-i\bm{q}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})|e^{i\theta_{\bm{k}\alpha}(\bm{q})}|u_{\beta}(\bm{k}+\bm{q})u_{\beta}^{*}(\bm{k})|e^{-i\theta_{\bm{k}\beta}(\bm{q})},$ (41)
This may lead to an unphysical contribution. In the expression before the projection, the quantity does not depend on the phase factor $\theta_{\alpha}(\bm{q})$. When the phase factor vanishes, the assumption in Eq.(Uniform pairing condition and minimum quantum metric) is valid. For it, we can shift the orbital positions by $\bm{\delta}_{\alpha}\rightarrow\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$, then
$\exp\left[-i(\theta_{\bm{k}\alpha}(\bm{q})-\theta_{\bm{k}\beta}(\bm{q}))\right]\rightarrow\exp\left[-i(\theta_{\bm{k}\alpha}(\bm{q})-\theta_{\bm{k}\beta}(\bm{q}))\right]e^{-i\bm{q}\cdot(\bm{x}_{\alpha}-\bm{x}_{\beta})},$ (42)
@@
$\frac{1}{N}\sum_{\bm{k}}\frac{d}{dq_{a}}\theta_{\bm{k}\alpha\beta}(\bm{q})|_{\bm{q}\rightarrow 0}=\bm{x}_{\alpha}^{a}-\bm{x}_{\beta}^{a}~{}.$ (43)
In this case, we will find that the quantum metric indeed becomes the minimal one. We can fix the position of the first orbital $\bm{\delta}_{1}=0$ and $\bm{x}_{1}=0$. The shift of other orbitals can be accordingly determined by Eq. (43). Furthermore, due to the dependence of $\theta_{\bm{k}\alpha\beta}(\bm{q})$ on the momentum $\bm{k}$, in general it is impossible to choose a set of orbital positions such that the second term in Eq. (Uniform pairing condition and minimum quantum metric) vanishes. This is different from the case of superfluid weight. In fact, the best we can do is to reach the minimal quantum metric in Eq. (Uniform pairing condition and minimum quantum metric). In Ref.Huhtinen et al. (2022), it was shown that the resulting $\bm{\delta}_{\alpha}+\bm{x}_{\alpha}$ for the minimal quantum metric is unique and independent of the initial positions $\bm{\delta}_{\alpha}$. Therefore, we can reach a conclusion that the coherence length $\xi$ for a superconductor is controlled by the minimal quantum metric under the uniform pairing condition.
One may find that the coherence length obtained under the uniform pairing condition is slightly larger than the value by directly evaluating the $\mathcal{M}(\bm{q})$ in Eq. (Uniform pairing condition and minimum quantum metric). There are two origins of the errors. Suppose that we start with the orbital positions corresponding to the minimal quantum metric. The first type of the error comes from the factor $\mathcal{M}(0)$ in Eq. (39), and the error can be evaluated from the Chebyshev’s sum inequality,
$\frac{1}{N}\sum_{\bm{k}\alpha}|u_{\alpha}(\bm{k})|^{4}-\frac{1}{NN_{orb}}\sum_{\bm{k}}(\sum_{\alpha}|u_{\alpha}(\bm{k})|^{2})^{2}=\frac{1}{NN_{orb}}\sum_{\bm{k},\alpha\beta}(|u_{\alpha}(\bm{k})|^{2}-|u_{\beta}(\bm{k})|^{2})^{2}\geq 0.$ (44)
@@
$\sigma^{2}_{ab}=\frac{1}{2N}\sum_{\bm{k}\alpha\beta}(\partial_{a}\theta_{\bm{k}\alpha\beta}(0)-\overline{\partial_{a}\theta_{\alpha\beta}(0)})(\partial_{b}\theta_{\bm{k}\alpha\beta}(0)-\overline{\partial_{b}\theta_{\alpha\beta}(0)}),$ (45)
@@
Supplementary Note 3: Coherence length for a weakly dispersive band
Derivation on the total coherence length
@@
$|\Lambda(\bm{k}+\bm{q},\bm{k})|^{2}$ $=1-\sum_{ab}\mathcal{G}_{ab}(\bm{k})q_{a}q_{b}+\mathcal{O}(q^{3}),$ (46)
@@
$\mathcal{G}_{ab}(\bm{k})=\mathrm{Re}\left[\langle\partial_{a}u(\bm{k})|\partial_{b}u(\bm{k})\rangle-\langle\partial_{a}u(\bm{k})|u(\bm{k})\rangle\langle u(\bm{k})|\partial_{b}u(\bm{k})\rangle\right],$ (47)
Instead of Eq. (Proof of the independence of coherence length from orbital positions) which only involves the normal Gor’kov’s Green function, we can consider a more compacted form by combining the Cooper-pair correlator and the density-density correlator,
$\mathcal{C}(\bm{r},\bm{r}^{\prime})$ $=\sum_{\alpha}\sum_{\mathcal{O}\in\mathcal{S}_{\mathrm{op}}}\langle\mathcal{O}(\bm{r})\mathcal{O}^{\dagger}(\bm{r}^{\prime})\rangle_{c}$
$\rightarrow\frac{T}{N}\sum_{\bm{k},\bm{q}}|\Lambda(\bm{k}+\bm{q},\bm{k})|^{2}\mathrm{Tr}G(i\omega_{n},\bm{k}+\bm{q})G(-i\omega_{n},-\bm{k})$ (48)
where $\mathcal{S}_{\mathrm{op}}=\{a_{\alpha\downarrow}a_{\alpha\uparrow},a^{\dagger}_{\alpha\uparrow}a^{\dagger}_{\alpha\downarrow},a^{\dagger}_{\alpha\downarrow}a_{\alpha\downarrow},a^{\dagger}_{\alpha\uparrow}a_{\alpha\uparrow}\}$ and for the notation simplification, we ignore the overall $1/N_{\mathrm{orb}}$ factor in this section. We make this treatment is to evaluate the coherence length analytically. In the last line, we have introduced the projection. Since the Green function in Eq. (Proof of the independence of coherence length from orbital positions) is independent on the momentum in the flatband limit, we can obtain the tendency
@@
$=\frac{\tanh\left(\frac{\beta\sqrt{\Delta^{2}+\mu^{2}}}{2}\right)}{\sqrt{\Delta^{2}+\mu^{2}}}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\frac{1}{1+\sum_{ab}\overline{\mathcal{G}}_{ab}q_{a}q_{b}},$ (49)
@@
$=$ $T\sum_{n}\frac{2}{\omega_{n}^{2}+\mu{}^{2}+\Delta^{2}}=\frac{\tanh\left(\frac{\beta\sqrt{\Delta^{2}+\mu^{2}}}{2}\right)}{\sqrt{\Delta^{2}+\mu^{2}}},$ (50)
@@
$=\frac{\tanh\left(\frac{\beta}{2}\varepsilon\right)}{\varepsilon\sqrt{\det\bar{\mathcal{G}}_{ab}}}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}}{q^{2}+\xi^{-2}},$ (51)
@@
$\xi=(\det\overline{\mathcal{G}}_{ab})^{\frac{1}{4}}.$ (52)
@@
$\mathcal{C}(\bm{r},\bm{r}^{\prime})\equiv\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\mathcal{M}(\bm{q}),$ (53)
@@
$\mathcal{M}(\bm{q})=\mathcal{M}_{0}+\mathcal{M}_{\mathrm{qm}}(\bm{q})+\mathcal{M}_{\mathrm{con}}(\bm{q}).$ (55)
@@
$\equiv\mathcal{M}_{0}(\Delta,\mu),$ (56)
@@
$\nu=\int\frac{d^{2}k}{(2\pi)^{2}}\theta(\mu_{0}-\epsilon(\bm{k})),$ (57)
@@
$\rho_{0}(\epsilon_{F})=\frac{1}{W}.$ (58)
Then we can calculate $\mathcal{M}_{0}(\Delta,\mu)$ in Eq. (56) at zero temperature,
@@
$=\frac{1}{\sqrt{\Delta^{2}+\mu^{2}}}.$ (59)
@@
$=-q^{2}\mathcal{M}_{0}(\Delta,\mu)\sqrt{\det\overline{\mathcal{G}}_{ab}},$ (60)
@@
$\overline{\mathcal{G}}_{ab}=\frac{\frac{1}{N}\sum_{\bm{k}}\mathcal{G}_{ab}(\bm{k})\frac{1}{\varepsilon(\bm{k})}}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\mathcal{G}_{ab}(\bm{k})}{\varepsilon(\bm{k})}}{\mathcal{M}_{0}(\Delta,\mu)},$ (61)
@@
$=-\frac{T}{N}\sum_{n\bm{k}}\frac{2\bm{v}_{F}\cdot\bm{q}[\epsilon(\bm{k})-\mu+\bm{v}_{F}\cdot\bm{q}]}{[\Delta^{2}+(\epsilon(\bm{k})-\mu)^{2}+\omega_{n}^{2}][\Delta^{2}+(\epsilon(\bm{k})-\mu+\bm{v}_{F}\cdot\bm{q})^{2}+\omega_{n}^{2}]},$ (62)